# Non-commutative Iwasawa theory of abelian varieties over global function fields

### v1

@@ 9. Computations of the Generalised Euler Characteristics
In order to compute the Generalised Euler characteristics, we impose in this section Hypothesis $(\mathbf{H})$ on the abelian variety $A$ defined over a function field $F$, which we now recall:
Hypothesis $(\mathbf{H})$. $H^{i}(H,A_{p^{\infty}}(F_{\infty}))$ is finite for all $i\geq 0$.
@@
Let $A$ be an abelian variety defined over a global field $F$ with $\mathrm{char}F\neq p$ and let $F_{\infty}$ be an admissible $p$-adic Lie extension of $F$. Suppose that Hypothesis $(\mathbf{H})$ holds. Then ${\mathrm{Sel}}(A/F_{\infty})$ has finite generalised $G$-Euler characteristic if and only if ${\mathrm{Sel}}(A/F^{\rm cyc})$ has finite generalised $\Gamma$-Euler characteristics. In the event of such, we have the equality
@@
where, for every $v\in S^{\prime}$, $w$ is a fixed prime of $F^{\mathrm{cyc}}$ above $v$.
@@
[…]
From the proof of Lemma 5.4, we see that the $p$-cohomological dimension of the decomposition group of $H$ at every prime of $F_{\infty}$ is at most one. It follows from this that the map $\gamma$ is surjective which in turn implies the same for $\phi$. The proof of the same lemma also gives $H^{i}(H,J_{v}(A/F_{\infty}))=0$ for every $i\geq 1$. Applying the snake lemma to the diagram (9.1), we obtain the following exact sequence
(9.2) $0\to\ker\alpha\to H^{1}(H,A_{p^{\infty}}(F_{\infty}))\to\ker\gamma\to\mathrm{coker}~{}\!\alpha\to H^{2}(H,A_{p^{\infty}}(F_{\infty}))\to 0$
and
$H^{i}(H,{\mathrm{Sel}}(A/F_{\infty}))\simeq H^{i+2}(H,A_{p^{\infty}}(F_{\infty}))$
for every $i\geq 1$. In view of this latter identification and Hypothesis $(\mathbf{H})$, we see that $H^{i}(H,{\mathrm{Sel}}(A/F_{\infty}))$ is finite for ech $i\geq 1$. It then follows from Lemma 8.3 that ${\mathrm{Sel}}(A/F_{\infty})$ has finite generalised $G$-Euler characteristic if and only if the map
@@
has finite kernel and cokernel. Since $\mathrm{coker}~{}\!h=H^{1}(H,{\mathrm{Sel}}(A/F_{\infty}))^{\Gamma}\simeq H^{3}(H,A_{p^{\infty}}(F_{\infty}))^{\Gamma}$ is also finite, we see that the map $d_{0}$ has finite kernel and cokernel if and only if $g$ does. The latter is precisely equivalent to saying that ${\mathrm{Sel}}(A/F_{\infty})^{H}$ has finite generalised $\Gamma$-Euler characteristics.
@@
(9.3) $0\to\ker\alpha\to{\mathrm{Sel}}(A/F^{\rm cyc})\to{\mathrm{Sel}}(A/F_{\infty})^{H}\to\mathrm{coker}~{}\!\alpha\to 0$
Since $\ker\alpha\subseteq H^{1}(H,A_{p^{\infty}}(F_{\infty}))$, we see that $\ker\alpha$ is finite. We claim that $\mathrm{coker}~{}\!\alpha$ has finite $\Gamma$-Euler characteristic. Suppose for now the claim holds. Then it follows from Lemma 8.5 and (9.3) that ${\mathrm{Sel}}(A/F_{\infty})^{H}$ has finite generalised $\Gamma$-Euler characteristics if and only if ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ has finite generalised $\Gamma$-Euler characteristics, establishing the first assertion of the theorem. It therefore remains to prove our claim. Taking Hypothesis $(\mathbf{H})$ into account, the exact sequence (9.2) reduces our task to showing that $\ker\gamma$ has finite $\Gamma$-Euler characteristic. For this, we note that as observed in the proof of Theorem 7.7, there is a surjection
with $H^{1}(F_{v},A_{p^{\infty}})$ being finite. In particular, $J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}$ is finite, and this in turn implies that $J_{v}(A/F^{\mathrm{cyc}})$ has a finite $\Gamma$-Euler characteristic.
From now on, we suppose that both generalised Euler characteristics in question are well-defined. Then we may apply Theorem 5.5 and Proposition 8.4 to conclude that
@@
Therefore, the final assertion of the theorem will follow once we can show that
@@
for each $v$. The restriction map yields a short exact sequence
@@
where Tate duality tells us that $H^{1}(F_{v},A_{p^{\infty}})$ is isomorphic to the Pontryagin dual of $A^{*}_{p^{\infty}}(F_{v})$. Thus, we have
(9.6) #$\#J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}=\frac{\#A^{*}_{p^{\infty}}(F_{v})}{\#H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{{\mathrm{cyc}}}_{w}))}.$
On the other hand, we have
(9.7) $H^{1}\left(\Gamma,J_{v}(A/F^{\mathrm{cyc}})\right)\simeq H^{1}(\Gamma_{w},H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}}))$
by Shapiro’s lemma. Now, consider the following spectral sequence
@@
Since $\mathrm{cd}_{p}~{}\Gamma=1$, $H^{2}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}})=0$ (cf. [28, Theorem 7.1.8(i)]) and $H^{2}(F_{v},A_{p^{\infty}})=0$ (see [11, Lemma 1.12]), the spectral sequence degenerates yielding
(9.8) $H^{1}(\Gamma_{w},H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}}))=0.$
Combining (9.6), (9.7) and (9.8), we obtain (9.5), as required. This completes the proof of our theorem. ∎
@@
where $L_{v}(E,1)$ denotes the Euler factor at $v$ of the complex $L$-function of $E$ over $F$, and $|\cdot|_{p}$ is the $p$-adic valuation of ${\mathbb{Q}}$, normalised so that $|p|_{p}=p^{-1}$. This recovers the formula of Zerbes in the number field setting and a formula of Sechi using the standard Euler characteristics in the function field setting.
@@
Retain the settings of Theorem 9.1. Suppose further that Greenberg’s conjecture semi-simplicity conjecture 6.2 holds for $A/F$. Set $s_{A}=\mathrm{corank}_{{\mathbb{Z}}_{p}}{\mathrm{Sel}}(A/F)$. We then have
$\frac{1}{T^{s_{A}}}\Phi_{\mathrm{reg}_{F}}(\xi_{A})\Big{|}_{T=0}=\chi(\Gamma,{\mathrm{Sel}}(A/F^{\mathrm{cyc}}))\prod_{v\in S^{\prime}}\frac{\#A^{*}_{p^{\infty}}(F_{v})}{\#H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{\mathrm{cyc}}_{w}))}.$
where, for every $v\in S^{\prime}$, $w$ is a fixed prime of $F^{\mathrm{cyc}}$ above $v$.

### v2

@@ 9. Computations of the Generalised Euler Characteristics
To compute the generalised Euler characteristics, we impose the following assumption $(\mathbf{H})$ on the abelian variety $A$ defined over a function field $F$, which we now recall:
$\mathbf{(H)}:\quad H^{i}(H,A_{p^{\infty}}(F_{\infty}))\text{ is finite for all }i\geq 0.$
@@
Let $A$ be an abelian variety defined over a global field $F$ with $\mathrm{char}F\neq p$, and let $F_{\infty}$ be an admissible $p$-adic Lie extension of $F$. Suppose that assumption $(\mathbf{H})$ holds. Then ${\mathrm{Sel}}(A/F_{\infty})$ has a finite generalised $G$-Euler characteristic if and only if ${\mathrm{Sel}}(A/F^{\rm cyc})$ has a finite generalised $\Gamma$-Euler characteristic. In that case, we have the following equality
@@
where, for each $v\in S^{\prime}$, $w$ is a fixed prime of $F^{\mathrm{cyc}}$ lying above $v$, and $\Gamma_{w}$ denotes the decomposition subgroup of $\Gamma$ at $w$.
@@
(9.1) ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Sel}}(A/F^{\rm cyc})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${H^{1}(G_{S}(F^{\rm cyc}),A_{p^{\infty}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta}$${\bigoplus_{v\in S}J_{v}(A/F^{\rm cyc})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Sel}}(A/F_{\infty})^{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}})^{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${\bigoplus_{v\in S}J_{v}(A/F_{\infty})^{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(H,{\mathrm{Sel}}(A/F_{\infty}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\cdots.}$
The rows are exact. From the proof of Lemma 5.4, we know that the $p$-cohomological dimension of the decomposition group of $H$ at every prime of $F_{\infty}$ is at most one. It follows that the map $\gamma$ is surjective. By applying the Snake lemma, it follows that the map $\phi$ is also surjective. Applying the snake lemma to the diagram (9.1), we obtain the following exact sequence
(9.2) $0\to\ker\alpha\to H^{1}(H,A_{p^{\infty}}(F_{\infty}))\to\ker\gamma\to\mathrm{coker}~\!\alpha\to H^{2}(H,A_{p^{\infty}}(F_{\infty}))\to 0.$
From Lemma 5.4, we have $H^{i}(H,J_{v}(A/F_{\infty}))=0$ for every $i\geq 1$. Noting that $\phi$ is surjective and using the long exact sequence in HH–cohomology associated to the short exact sequence (4.2), we obtain an isomorphism
$H^{i}(H,{\mathrm{Sel}}(A/F_{\infty}))\simeq H^{i}(H,H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}}))\qquad\textrm{for all $i\geq 1$.}$
Consider the Hochschild–Serre spectral sequence
$H^{p}(H,H^{q}(G_{S}(F_{\infty}),A_{p^{\infty}}))\Rightarrow H^{p+q}(G_{S}(F^{\rm cyc}),A_{p^{\infty}}).$
From Lemma 3.2, for every $i\geq 1$, we have
$H^{i}(H,H^{1}(G_{S}(F_{\infty}),A_{p^{\infty}}))\simeq H^{i+2}(H,A_{p^{\infty}}(F_{\infty})).$
Thus, for all $i\geq 1$,
$H^{i}(H,{\mathrm{Sel}}(A/F_{\infty}))\simeq H^{i+2}(H,A_{p^{\infty}}(F_{\infty})).$
By assumption $(\mathbf{H})$, $H^{i}(H,{\mathrm{Sel}}(A/F_{\infty}))$ is finite for all $i\geq 1$. Consider now the following exact sequence
$0\to H^{1}(\Gamma,H^{i}(H,{\mathrm{Sel}}(A/F_{\infty})))\to H^{i+1}(G,{\mathrm{Sel}}(A/F_{\infty}))\to H^{0}(\Gamma,H^{i+1}(H,{\mathrm{Sel}}(A/F_{\infty})))\to 0.$
From Lemma 8.3, ${\mathrm{Sel}}(A/F_{\infty})$ has a finite generalised $G$-Euler characteristic if and only if the map
@@
has a finite kernel and cokernel. Since $\mathrm{coker}~\!h=H^{1}(H,{\mathrm{Sel}}(A/F_{\infty}))^{\Gamma}\simeq H^{3}(H,A_{p^{\infty}}(F_{\infty}))^{\Gamma}$ is also finite, it follows that $d^{0}$ has a finite kernel and cokernel if and only if $g$ does. The latter is precisely equivalent to saying that ${\mathrm{Sel}}(A/F_{\infty})^{H}$ has a finite generalised $\Gamma$-Euler characteristic.
@@
(9.3) $0\to\ker\alpha\to{\mathrm{Sel}}(A/F^{\rm cyc})\to{\mathrm{Sel}}(A/F_{\infty})^{H}\to\mathrm{coker}~\!\alpha\to 0.$
Since $\ker\alpha\subseteq H^{1}(H,A_{p^{\infty}}(F_{\infty}))$, we know $\ker\alpha$ is finite. We claim that $\mathrm{coker}~\!\alpha$ has a finite $\Gamma$-Euler characteristic. Assuming the claim, it follows from Lemma 8.5 and (9.3) that ${\mathrm{Sel}}(A/F_{\infty})^{H}$ has a finite generalised $\Gamma$-Euler characteristic if and only if ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ does. This establishes the first assertion of the theorem. To prove the claim, note that by (9.2), it suffices to show that $\ker\gamma$ has a finite $\Gamma$-Euler characteristic. As observed in the proof of Theorem 7.7, there is a surjection
$H^{1}(F_{v},A_{p^{\infty}})\twoheadrightarrow J_{v}(A/F^{\mathrm{cyc}})^{\Gamma},$
and since $H^{1}(F_{v},A_{p^{\infty}})$ is finite, so is $J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}$. Thus, we see that $({\rm ker}\gamma)^{\Gamma}$ is also finite. By [12, Proposition A.1.7], it follows that ${\rm ker}(\gamma)$ has a finite $\Gamma$-Euler characteristic.
Now suppose that both generalised Euler characteristics in question are well-defined. Then by Theorem 5.5 and Proposition 8.4, we obtain
@@
Thus, the final assertion of the theorem will follow once we show that
@@
for each $v\in S^{\prime}$. The restriction map gives the following short exact sequence
@@
where, using Shapiro’s lemma, we identify
$J_{v}(A/F^{\rm cyc})^{\Gamma}\simeq H^{0}(\Gamma_{w},H^{1}(F^{\rm cyc}_{w},A_{p^{\infty}})).$
By Tate local duality, $H^{1}(F_{v},A_{p^{\infty}})$ is dual to $A^{*}_{p^{\infty}}(F_{v})$. Therefore,
(9.6) $J_{v}(A/F^{\mathrm{cyc}})^{\Gamma}=\frac{\#A^{*}_{p^{\infty}}(F_{v})}{\#H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{{\mathrm{cyc}}}_{w}))}.$
On the other hand, by Shapiro’s lemma, we have
(9.7) $H^{1}\left(\Gamma,J_{v}(A/F^{\mathrm{cyc}})\right)\simeq H^{1}(\Gamma_{w},H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}})).$
Now consider the following spectral sequence
@@
Since $\mathrm{cd}_{p}~\Gamma_{w}=1$ and $H^{2}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}})=0$ (cf. [35, Theorem 7.1.8(i)]), the spectral sequence degenerates, yielding an isomorphism
(9.8) $H^{1}(\Gamma_{w},H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}}))\cong H^{2}(F_{v},A_{p^{\infty}}).$
By Tate local duality, $H^{2}(F_{v},A_{p^{\infty}})=0$ (see [12, Lemma 1.12]). Therefore, combining (9.6), (9.7) and (9.8), we obtain the identity (9.5), as required. This completes the proof of the theorem. ∎
@@
where $L_{v}(E,1)$ denotes the Euler factor at $v$ of the complex $L$-function of $E$ over $F$. This recovers a formula of Zerbes in the number field setting and a corresponding result of Sechi in the function field setting using standard Euler characteristics.
In [24, Theorem 3.2.6], Lai–Longhi–Tan–Trihan computed the generalised $\Gamma$-Euler characteristic
$\chi(\Gamma,{\mathrm{Sel}}(A/F^{\mathrm{cyc}}))$
in terms of the quantities appearing in the Birch–Swinnerton-Dyer conjecture. Combining their result with ours, we deduce that the generalised $G$-Euler characteristic $\chi(G,{\mathrm{Sel}}(A/F_{\infty}))$ can likewise be expressed in terms of invariants arising in the Birch–Swinnerton-Dyer conjecture. We conclude with the following corollary, which illustrates how the characteristic element fits into this framework.
@@
Retain the setting of Theorem 9.1. Suppose further that Greenberg’s semi-simplicity conjecture 6.2 holds for $A/F$. Set $s_{A}=\mathrm{corank}_{{\mathbb{Z}}_{p}}{\mathrm{Sel}}(A/F)$. Then
$\frac{1}{T^{s_{A}}}\Phi_{\mathrm{reg}_{F}}(\xi_{A})\Big|_{T=0}=\chi(\Gamma,{\mathrm{Sel}}(A/F^{\mathrm{cyc}}))\prod_{v\in S^{\prime}}\frac{\#A^{*}_{p^{\infty}}(F_{v})}{\#H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{\mathrm{cyc}}_{w}))},$
where, for each $v\in S^{\prime}$, $w$ is a fixed prime of $F^{\mathrm{cyc}}$ lying above $v$, and $\Gamma_{w}$ is the corresponding decomposition subgroup.
@@
Acknowledgements
The authors thank Ashay Burungale, Pierre Colmez, Mahesh Kakde, Antonio Lei, Katharina Müller, Anwesh Ray, Jishnu Ray, Gianluigi Sechi, Ramdorai Sujatha, Fabien Trihan, Maria Valentino and Sarah Zerbes for their interest and comments. They are very grateful to the anonymous referee for valuable suggestions. This paper arose from the conference “Special Values of $L$-functions” held at Tsinghua Sanya International Mathematics Forum, and the authors thank Ashay Burungale, Emmanuel Lecouturier and Xin Wan for organising the conference that enabled the authors’ collaboration. Yukako Kezuka also thanks Gianluigi Sechi for sharing his PhD thesis, which inspired their collaboration. Yukako Kezuka was partially supported by the ANR project Coloss, the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101026826, and JSPS KAKENHI Grant JP25K17227.
Conflict of interest. On behalf of all authors, the corresponding author states that there is no conflict of interest.